# A Human-Centered Approach to Identifying Promises, Risks, & Challenges of Text-to-Image Generative AI in Radiology

### v1

@@ Introduction.
Impressively, the model was able to generate pathologies that were not tested in the original paper (Xu et al. 2024). However, there are four primary technical and AI safety challenges that arose while participants were generating and exploring CT scans: (1) the resolution quality of the CT Scans, (2) the approach to validation and hallucinations, (3) generating contrast-enhanced CTs, and (4) technical expectations of the model generation.
Five participants commented on the resolution of the generated CT scans being of lower quality, even though the model we used was designed to develop high-resolution CTs. While looking through their generated CT scan, R7 found it difficult to see if the pathology they asked for was there “because of the resolution”. They described the experience as if they were looking at a CT scan on a really bad computer instead of the expensive computers that the radiologists use, while R9 described it looking “like they moved during” the CT scan. R6 similarly asked during their session if they had a bad internet connection or if it was just the generation quality. However, R3 clarified that a radiologist may not necessarily require a high-resolution CT scan as long as it offers the required information to support the goal of the generation: “It’s not necessary to fool me into thinking it’s a real CT. What you want it to do is provide a sufficient example that I would recognize this on a real CT”.
Second, several participants encountered flaws in the generated output as they were prompting the model on image findings that were not tested for by the quantitative model evaluations. For example, R3 spent time closely auditing the outputs, commenting that the model has “…thrown all sorts of pathology onto this patient…” including “…metastatic bone disease…and a plural recess down here that doesn’t connect to anything”. R4 and R9 similarly identified missing or injected pathologies in their generations. A vast majority of the ML literature refers to this type of model behavior as hallucinations, which come with domain-specific risks. As with any black box AI, it is important that users can see which part of the prompts are associated with which regions of the generation. This becomes especially critical and challenging with a 3D-generated output where multiple pathologies are combined within a single prompt. This task, however, is computationally expensive and underexplored for text-to-image GenAI.
@@
Lastly, despite the GenAI’s potential value, R3 and R9 acknowledge that the generation time delay is substantial (33~{}33 minutes), especially if comparing it to searching for examples on Radiopaedia. R9 said that as residents, they have “…limited time while you’re reading images to do research.” Beyond generation time, several participants commented on the GenAI’s limitations compared to content provided by existing resources, such as Radiopaedia. For example, R5 contrasts the model & interface to how “Radiopaedia also has some additional notes [about the image finding] which can be helpful for differential diagnosis”. These resources allow them to parse detailed descriptions, aiding in their conclusion about a diagnosis, whereas the GenAI is limited to taking a single prompt at a time to generate a CT scan and does not provide any detailed description in return.
@@
The development of medical T2I that addresses stakeholders’ needs and concerns goes beyond solving technical challenges. We must also consider the complex interplay between user expectations and domain-specific implications. Our findings revealed two domain-specific risks that need to be considered to ensure that T2I GenAI meets the needs of the key stakeholders and aligns with complex workflows: (1) mitigating confirmation bias and the distillation of incorrect information and (2) supporting the need to continuously be exposed to real medical images.
First, some participants were unsure if what they were looking at was what the AI generated for that pathology or not (R2, R6, R7, R9). Several participants would say something like “Maybe this is what they wanted” (R2) or “I potentially see that dissection” (R7) while reviewing their generations. This type of behavior with AI, masked as confirmation bias, can be detrimental to a trainees learning progress. R9 highlighted potential consequences: “If you’re learning and then you get like false information, uh oh, you know, that’s confusing. And you don’t know if you’re wrong or [the] software is wrong”. In the use case of practice exams, these unprompted pathologies could distract more than support the student (R3). Ultimately, it is possible for students generating CT scans who are unaware of the additional, missing, or misrepresented pathologies could unintentionally learn the wrong things. For example, after R7’s generated a CT scan for aortic dissection, they describe a potential issue related to confirmation bias: “Am I seeing what I’m actually seeing? Or am I seeing what I want to see?” To this point, R7 emphasizes that research would need to “…make sure that this is something that’s not going to…indirectly, [give] poor quality education.
Second, despite the participants acknowledging that the generated CT scans were of lower resolution, they were surprised by the fact that they were looking at a generated CT scan. R7 remarked, “If you hadn’t told me I wouldn’t have known [that the CT was AI generated]”. F2 reacted similarly, saying “Oh wow that’s really good because I couldn’t tell which one is from which”. R5 stated “With the right prompt, the quality of the output is so good that I cannot distinguish that this is AI, if I did not create the prompt myself”. This in itself raises potentially consequential implications as R5 states that “Generated images can be useful for experienced doctors who have seen thousands of images, but residents need to develop experience with real patient cases and images”. From a lecturer’s perspective, the students want “…to see real examples with validated pathology” (R3). This underscores the ongoing theme that radiologists and those in the broader medical field need to continue seeing a plethora of real CT scans throughout their career, since the ones they need to interpret that matter are from real patients.
@@ Discussion
Building advanced models, such as the text-to-CT Scan GenAI, often occurs without a deep understanding of key stakeholders’ needs and potential safety risks. This lack of clarity raises important questions about misuse, misinformation, and the potential erosion of expert skills, particularly in high-stakes environments such as medicine (Karabacak et al. 2023; Ooi et al. 2023). Our work builds on and formalizes the growing body of research highlighting the potential of text-to-image GenAI beyond traditional applications. Our work also extends ongoing RAI research by identifying technical challenges and domain-specific risks of generating synthetic medical images. Below, we discuss the implications associated with the use of text-to-image GenAI in medical education, training, and practice.
Implications for Evaluating Text-to-Image GenAI
Human-Centered T2I Evaluations with Stakeholders. We found that each participant in our study was interested in exploring and evaluating a different capability of the model, which uncovered numerous technical challenges and domain-specific risks. Despite including quality feedback questions for each generation in our user interface, participants did not utilize them as they were not aligned with what they really wanted to audit: the ability to generate contrast-enhanced CT scans, the resolution of the images (similar to Brisco, Hay, and Dhami (2023)’s observation), or the misrepresentation of image findings.
We also learned that CT scan impressions can be subjective, making it important to be concerned about the demographics of the radiologists who wrote the reports for the training data. Using reports by trainees or unverified by multiple radiologists could introduce inconsistencies that can impact the reliability of the model. All of these challenges are difficult to quantitatively capture; however, conducting exploratory human-centered AI evaluations to simultaneously elicit this will be useful to the advancement of RAI.
@@
Our participants emphasized the role that online resources and pure experience play when faced with challenging cases and professional growth, as similarly observed by Yildirim et al. (2024) and Rassie (2017). El-Ali et al. (2019) and  Quaia (2023) have observed an increase in students’ and trainees’ performance when provided access to radiology-specific repositories, emphasizing the importance of supplemental resources to traditional lecture materials. Extending these insights, we learned that getting exposure to variations can play an important role in competency within medical imaging-based professions. Generating variations as a design has similarly been explored by (Gero et al. 2024), highlighting the value that variation plays in sensemaking. However, our design probe did not easily enable the generation and visualization of variations for a single prompt. This was also difficult due to the generation time for a single CT scan. As Weisz et al. (2024) suggest, it is important to design GenAI applications to enable variability and expose differences from a prompt. While this generates variability aimed at meeting the user’s needs, we also see how this can expedite education by showing several CT scans with the same pathology in several variations. This would help prepare students to practice identifying serious conditions before dealing with real patients, a benefit that del Cerro and Esteve (2023) similarly mention.
Our participants also stressed the importance of being able to understand what real human anatomy is like by looking at thousands of real CT scans. This opens the door for RAI work to explore various design opportunities for domain-specific T2I models, such as inpainting capabilities. Furthermore, RAI can explore the human-centered needs around explicitly identifying which case in the viewing window is generated and which is the real one to avoid misuse and abuse of the generative capabilities (del Cerro and Esteve 2023).
@@
Safeguard Against Hallucinations & Confirmation Bias. As we have seen through our own experiences and previous work, hallucinations are prevalent in GenAI systems. This is especially concerning because primarily the more experienced participants were commenting on misrepresentations of the prompt in their generated CT scan. This motivates future works to explore ways to express hallucinations to users, especially to novice users. Along with hallucinations, confirmation bias can emerge when working with AI (Rosbach et al. 2025), especially among students and junior trainees. While it may be hard to avoid generating hallucinations, we can design mechanisms to safeguard end-users from confirmation bias during their learning experience. Designing mechanisms to minimize confirmation bias is especially important in an educational environment because recent work has shown how people can start reasoning incorrectly when given misleading explanations generated by AI (Spitzer et al. 2024). These designs are especially important to avoid deskilling, slowing professional growth, and minimizing over-reliance. One potential solution that we see is by developing visual explanations that associate concepts from the prompt to regions of the generated image, such as the technical solutions proposed by Chefer et al. (2023) and Evirgen, Wang, and Chen (2024).
Preventing Misuse and Malpractice. Koohi-Moghadam and Bae (2023) suggests the potential for AI-generated medical images to fabricate clinical results or patients’ conditions, which can lead to inaccurate conclusions or mistreating patients. The misuse and malpractice scenarios could go on, but it is more important to discuss practices to prevent such scenarios from becoming prevalent. This especially becomes important when these models increase competency by being able to do tasks such as inpainting original patient CT scans. Horvitz (2022) proposes watermarking as a potential solution, being pivotal to ensuring transparency. Ultimately, there needs to be a clear indicator between a generated CT and a patient’s real CT.
@@
Similarly, we did not perform quantitative evaluations of the accuracy of the generated CT scans from the radiologists’ perspective during these sessions. However, future work can consider including such model auditing activities within contextually accurate tools to optimize how medical stakeholders can best give feedback on the output quality in such a complex environment of 3D images. Conducting these audits and evaluations within familiar medical imaging interfaces, rather than generic survey platforms, is critical for maintaining external validity, as radiologists rely on specialized functionalities not typically available.
@@ Conclusion
Medical text-to-image GenAI capabilities are increasingly being investigated as models become increasingly competent. Given the lack of human-centered research involving key stakeholders on the promises, risks, and challenges of medical text-to-image GenAI, our work seeks to bridge this gap by conducting exploratory model prompting sessions to elicit the promises, risks, and challenges of generating synthetic medical imagery. Our study highlights the opportunities of text-to-image GenAI in medical education, training, and practice, such as empowering students to learn in low-stakes environments and augmenting radiologists when faced with challenging cases. These promising opportunities expose technical challenges and domain-specific risks of T2I models that the RAI community can consider. Overall, our work represents the importance of evaluating T2I models from a human-centered lens, taking domain-specific needs and challenges into consideration. We hope this work encourages RAI researchers to further explore the risks and challenges of domain-specific T2I models.

### v2

@@ Introduction.
Impressively, the model was able to generate pathologies that were not tested in the original study (Xu et al. 2024). However, there are four primary technical and AI safety challenges that arose while participants were generating and exploring CT scans: (1) the resolution quality of the CT Scans, (2) the approach to validation and hallucinations, (3) generating contrast-enhanced CTs, and (4) technical expectations of the model generation.
Five participants commented on the resolution of the generated CT scans being of lower quality, even though the model we used was designed to develop high-resolution CTs. While looking through their generated CT scan, R7 found it difficult to see if the pathology they asked for was there “because of the resolution”. They described the experience as if they were looking at a CT scan on a really bad computer instead of the expensive computers that the radiologists use, while R9 described it looking “like they moved during” the CT scan. R6 similarly asked during their session if they had a bad internet connection or if it was just the generation quality. However, R3 clarified that a radiologist may not necessarily require a high-resolution CT scan as long as it offers the required information to support the goal of the generation: “It’s not necessary to fool me into thinking it’s a real CT. What you want it to do is provide a sufficient example that I would recognize this on a real CT.”
Second, several participants encountered flaws in the generated output as they were prompting the model on image findings that were not tested for by the quantitative model evaluations. For example, R3 spent time closely auditing the outputs, commenting that the model has “…thrown all sorts of pathology onto this patient…” including “…metastatic bone disease…and a plural recess down here that doesn’t connect to anything.” R4 and R9 similarly identified missing or injected pathologies in their generations. A vast majority of the ML literature refers to this type of model behavior as hallucinations, which come with domain-specific risks. As with any black box AI, it is important that users can see which part of the prompts are associated with which regions of the generation. This becomes especially critical and challenging with a 3D-generated output where multiple pathologies are combined within a single prompt. This task, however, is computationally expensive and underexplored for text-to-image GenAI (Bradshaw et al. 2025).
@@
Lastly, despite the GenAI’s potential value, R3 and R9 acknowledge that the generation time delay is substantial (3~3 minutes), especially if comparing it to searching for examples on Radiopaedia. R9 said that as residents, they have “…limited time while you’re reading images to do research.” Beyond generation time, several participants commented on the GenAI’s limitations compared to content provided by existing resources, such as Radiopaedia. For example, R5 contrasts the model & interface to how “Radiopaedia also has some additional notes [about the image finding] which can be helpful for differential diagnosis.” These resources allow them to parse detailed descriptions, aiding in their conclusion about a diagnosis, whereas the GenAI is limited to taking a single prompt at a time to generate a CT scan and does not provide any detailed description in return.
@@
The development of medical T2I that addresses stakeholders’ needs and concerns goes beyond solving technical challenges. We must also consider the complex interplay between user expectations and domain-specific implications. Our findings revealed two domain-specific risks that need to be considered to ensure that T2I GenAI meets the needs of the key stakeholders and aligns with complex workflows: (1) mitigating confirmation bias and the distillation of incorrect information and (2) supporting continuous exposure to real medical images.
First, participants were unsure if and where the requested pathology from the prompt was generated (R2, R6, R7, R9). While reviewing the generations, they would say something like “Maybe this is what they [AI] wanted” (R2) or “I potentially see that dissection” (R7). This type of behavior with AI, masked as confirmation bias, can be detrimental to a trainee’s learning progress. R9 highlighted potential consequences: “If you’re learning and then you get like false information, uh oh, you know, that’s confusing. And you don’t know if you’re wrong or [the] software is wrong.” In the use case of practice exams, these unprompted pathologies could distract more than support the student (R3). Ultimately, it is possible for students generating CT scans who are unaware of the additional, missing, or misrepresented pathologies to unintentionally learn the wrong things. For example, after R7 generated a CT scan for aortic dissection, they describe a potential issue related to confirmation bias: “Am I seeing what I’m actually seeing? Or am I seeing what I want to see?” To this point, R7 emphasizes that research would need to “…make sure that this is something that’s not going to…indirectly, [give] poor quality education.
Second, despite the participants acknowledging that the generated CT scans were of lower resolution, they were surprised by the fact that they were looking at a generated CT scan. R7 remarked, “If you hadn’t told me I wouldn’t have known [that the CT was AI generated].” F2 reacted similarly, saying “Oh wow that’s really good because I couldn’t tell which one is from which.” R5 stated “With the right prompt, the quality of the output is so good that I cannot distinguish that this is AI, if I did not create the prompt myself.” This in itself raises potentially consequential implications as R5 states that “Generated images can be useful for experienced doctors who have seen thousands of images, but residents need to develop experience with real patient cases and images.” From a lecturer’s perspective, the students want “…to see real examples with validated pathology” (R3). This underscores the ongoing theme that radiologists and those in the broader medical field need to continue seeing a plethora of real CT scans throughout their careers, given that the ones they need to interpret are real patients.
@@ Discussion
Building advanced models, such as the text-to-CT Scan GenAI, often occurs without a deep understanding of key stakeholders’ needs and potential safety risks. This lack of clarity raises important questions about misuse, misinformation, and the potential erosion of expert skills, particularly in high-stakes environments such as medicine (Karabacak et al. 2023; Ooi et al. 2023). Below, we discuss the benefits of our study design and the implications associated with the use of text-to-image GenAI for radiology
Evaluating Domain-Specific Text-to-Image GenAI
Human-Centered Evaluations of T2I GenAI with Stakeholders. As an outcome of conducting exploratory model prompting with key stakeholders, we observed how each participant in our study was interested in exploring and evaluating a different capability of the model. Unlike the formative interviews, where users could not prompt the model, the exploratory model prompting study design allowed us to uncover numerous technical challenges and domain-specific risks. For example, it helped expose that the thumbs-up/down feedback mechanisms offered for each generated image in the interface did not address what participants really wanted to audit, which was the ability to generate contrast-enhanced CT scans, the resolution of the images (similar to Brisco, Hay, and Dhami (2023)’s observation), or the misrepresentation of image findings.
We also learned that CT scan impressions can be subjective, making it important to be concerned about the expertise of the radiologists who wrote the reports for the training data. Using reports by trainees or unverified by multiple radiologists could introduce inconsistencies that can impact the reliability of the model. All of these challenges are difficult to quantitatively capture; however, conducting exploratory human-centered AI evaluations to simultaneously elicit this will be useful to the advancement of RAI.
@@
We allude to RQ1 as we discuss the implications of T2I for medical education, training, and practice. Our participants emphasized the role that online resources and pure experience play when faced with challenging cases and professional growth, as similarly observed by Yildirim et al. (2024) and Rassie (2017). El-Ali et al. (2019) and  Quaia (2023) have observed an increase in students’ and trainees’ performance when provided access to radiology-specific repositories, emphasizing the importance of supplemental resources to traditional lecture materials. Extending these insights, five participants referenced how getting exposure to variations can play an important role in competency within medical imaging-based professions. Generating variations as a design has similarly been explored by (Gero et al. 2024), highlighting the value that variation plays in sensemaking. However, the design of our GenAI plugin did not easily enable the generation and visualization of variations for a single prompt. This was also difficult due to the generation time for a single CT scan. As Weisz et al. (2024) suggest, it is important to design GenAI applications to enable variability and expose differences from a prompt. While this generates variability aimed at meeting the user’s needs, we also see how this can expedite education by showing several CT scans with the same pathology in several variations. This would help prepare students to practice identifying serious conditions before dealing with real patients, a benefit that del Cerro and Esteve (2023) similarly mention.
Our participants also stressed the importance of being able to understand what real human anatomy is like by looking at thousands of real CT scans. This opens the door for RAI work to explore various design opportunities for domain-specific T2I models, such as inpainting capabilities. Furthermore, RAI can explore the human-centered needs around explicitly identifying which case in the viewing window is generated and which is real to avoid misuse and abuse of the GenAI capabilities (del Cerro and Esteve 2023).
@@
Safeguard Against Hallucinations & Confirmation Bias. As we have seen through our own experiences and previous work, hallucinations are prevalent in GenAI systems. This is especially concerning because primarily the more experienced participants were commenting on misrepresentations of the prompt in their generated CT scan. This motivates future works to explore ways to express hallucinations to users, especially to novice users. Along with hallucinations, some participants raised concerns about confirmation bias, which can emerge when working with AI (Rosbach et al. 2025), especially among students and junior trainees. While it may be hard to avoid generating hallucinations, we can design mechanisms to safeguard end-users from confirmation bias during their learning experience. Designing mechanisms to minimize confirmation bias is especially important in an educational environment because recent work has shown how people can start reasoning incorrectly when given misleading explanations generated by AI (Spitzer et al. 2024). These designs are especially important to avoid deskilling, slowing professional growth, and minimizing over-reliance. One potential solution that we see is by developing visual explanations that associate concepts from the prompt to regions of the generated image, such as the technical solutions proposed by Chefer et al. (2023)Bradshaw et al. (2025), and Evirgen, Wang, and Chen (2024).
Preventing Misuse and Malpractice. Although this was not an emerging theme from the interviews, it is necessary to anticipate potential misuse and malpractice of synthetic medical images. Koohi-Moghadam and Bae (2023) suggests the potential for AI-generated medical images to fabricate clinical results or patients’ conditions, which can lead to inaccurate conclusions or mistreating patients. The misuse and malpractice scenarios could go on, but it is more important to discuss practices to prevent such scenarios from becoming prevalent. This especially becomes important when these models increase competency by being able to do tasks such as inpainting original patient CT scans. Horvitz (2022) proposes watermarking as a potential solution to distinguish synthetic from real content, promoting transparency.
@@
Similarly, we did not perform quantitative evaluations of the accuracy of the generated CT scans from the radiologists’ perspective during these sessions. Future work should consider including quantitative model auditing activities with stakeholders using familiar medical interfaces and tools to maintain external validity and to ensure that evaluations appropriately target stakeholders’ concerns.
@@ Conclusion
Medical text-to-image GenAI capabilities are increasingly being investigated as models become increasingly competent. Given the lack of human-centered research involving key stakeholders on the promises, risks, and challenges of medical text-to-image GenAI, our work seeks to bridge this gap by conducting exploratory model prompting sessions with a text-to-CT Scan generative model. Our study highlights the opportunities of text-to-image GenAI in medical education, training, and practice, such as empowering students to learn in low-stakes environments and augmenting radiologists when faced with challenging cases. These promising opportunities expose technical challenges and domain-specific risks of domain-specific T2I models that the RAI community can consider, such as confirmation bias. Overall, our work represents the importance of evaluating T2I models from a human-centered lens, taking domain-specific needs and challenges into consideration. We hope this work encourages RAI researchers to further explore the risks and challenges of domain-specific T2I models.